# Untying the Reversal Curse via Bidirectional Language Model Editing

### v1

@@ Abstract
Recent studies have demonstrated that large language models (LLMs) store massive factual knowledge within their parameters. But existing LLMs are prone to hallucinate unintended text due to false or outdated knowledge. Since retraining LLMs is resource intensive, there has been a growing interest in the concept of model editing. Despite the emergence of benchmarks and approaches, these unidirectional editing and evaluation have failed to explore the reversal curse. Intuitively, if “The capital of France is” is edited to be a counterfact “London” within a model, then it should be able to naturally reason and recall the reverse fact, i.e., “London is the capital of” followed by “France” instead of “England”. In this paper, we study bidirectional language model editing, aiming to provide rigorous model editing evaluation to assess if edited LLMs can recall the editing knowledge bidirectionally. A new evaluation metric of reversibility is introduced and a benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed to evaluate the reversibility of edited models in recalling knowledge in the reverse direction of editing. We surprisingly observe that while current editing methods and LLMs can effectively recall editing facts in the direction of editing, they suffer serious deficiencies when evaluated in the reverse direction. To mitigate the reversal curse, a method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed. A set of editing objectives that incorporate bidirectional relationships between subject and object into the updated model weights are designed. Experiments show that BIRD improves the performance of four representative LLMs of different sizes via question answering and judgement.
@@
Figure 1: Comparison of existing unidirectional evaluation paradigms and the proposed paradigm that assesses model editing bidirectionally. Assessing whether edited models can recall the editing facts (a) via single-hop questions in the direction of editing; (b) via multi-hop questions in the direction of editing; (c) in the reverse direction of editing. $f_{\theta}$ / $f_{\theta_{e}}$ denotes the models before / after editing.
@@ 1 Introduction
Large language models (LLMs) have shown impressive capabilities in understanding and generating text, improving performance on a variety of natural language processing tasks such as code generation (Chen et al., 2021; Mitchell et al., 2022a), open-ended dialogue (Wu et al., 2023; Ouyang et al., 2022) and machine translation (Liu et al., 2022; Zhang et al., 2023a). However, existing LLMs inevitably exhibit hallucinations due to false or outdated knowledge stored in their parameters (Zhang et al., 2023b; Peng et al., 2023; Ji et al., 2023), which degrades the system performance and fails to meet user expectations in many real-world scenarios. Considering the huge computing resource consumption of retraining LLMs and echoing green deep learning (Schwartz et al., 2020; Xu et al., 2021), there has been considerable attention on the research regarding model editing (Sinitsin et al., 2020; Zhu et al., 2020; Cao et al., 2021; Meng et al., 2022, 2023; Dai et al., 2022; Mitchell et al., 2022b; Yao et al., 2023), which aims at modifying knowledge by updating model weights.
@@
Existing methods developed to modify the parameters of models can be generally categorized into two paradigms (Wang et al., 2023; Yao et al., 2023). (1) Meta-learning: these methods employ an editor hypernetwork to learn the adjustments ()($\triangle$)( △ ) for parameter updates in the base model (Cao et al., 2021; Mitchell et al., 2022a). (2) Locate-Then-Edit: these approaches locate knowledge neurons in LLMs that are positively correlated to a knowledge expression and then apply modifications to them accordingly (Dai et al., 2022; Meng et al., 2022, 2023). Currently, the effectiveness of an editing method is usually evaluated along three dimensions. (1) Efficacy is to validate whether the post-editing model can recall the exact editing fact. (2) Generalization is to verify whether the edited model can recall the editing fact under different paraphrase prompts. (3) Locality [^1] Also referred to as specificity in Meng et al. (2022, 2023). is to verify whether the output of the edited model for unrelated inputs remains unchanged after editing. Recently, researchers have made strides in model editing for LLMs in various editing tasks and settings. Both the ZsRE (Levy et al., 2017) and CounterFact (Meng et al., 2022) datasets utilize the aforementioned evaluation metrics to assess whether edited models can recall the newly injected facts as depicted in Figure 1(a), and remain unaffected for unrelated knowledge. To further assess the model editing capabilities, Yao et al. (2023) have taken a step forward to introduce an evaluation metric regarding portability, which is designed to gauge the effectiveness of model editing in transferring knowledge to related content. Besides, MQuAKE dataset (Zhong et al., 2023) is constructed to extend the assessment of generalization by measuring whether edited models correctly answer the multi-hop questions entailed by the editing facts, as shown in Figure 1(b). Despite the emergence of benchmarks and techniques, these editing and evaluation approaches operate under the unidirectional paradigm following only the direction being edited. The reversal curse (Berglund et al., 2023) has not yet been explored in model editing, and it remains unclear whether edited models can reason and recall the editing facts in the reverse direction of editing, as shown in Figure 1(c).
@@
In this paper, we study bidirectional language model editing, with the goal of conducting a comprehensive evaluation of model editing to determine if edited LLMs can accurately recall the editing knowledge bidirectionally. Among various kinds of learned beliefs such as logical, spatial, or numerical knowledge, this paper investigates factual knowledge within LLMs [^2] Model editing is also termed as knowledge editing.. In addition to the metrics of efficacy, generalization and locality that are commonly used in previous works, a new evaluation metric of reversibility is introduced to evaluate the effectiveness of edited models in recalling knowledge in the reverse direction of editing. To evaluate the reversibility of edited models, a benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed, where each piece of knowledge is represented in the form of a triple (subject, relation, object). When inverting the relation between subject and object, the taxonomy designed in Bordes et al. (2013) which categorizes relations between entities into four classes: one-to-one, one-to-many, many-to-one, many-to-many, is considered. Furthermore, two evaluation forms of question answering (QA) and judgment are adopted to characterize these symmetrical and asymmetrical relations.
To establish the performance comparison of existing model editing methods on bidirectional model editing, this paper presents meticulous experiments involving as many methods as possible such as FT (Zhu et al., 2020), MEND (Mitchell et al., 2022a), KN (Dai et al., 2022), ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023). Four representative LLMs of different sizes are selected as the base models: GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang & Komatsuzaki, 2021), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). We surprisingly observe that while current model editing methods and LLMs perform significantly well at recalling the editing facts in the editing direction, they suffer serious deficiencies when evaluated in the reverse direction. For instance, though a LLaMA-2 model edited by the state-of-the-art method ROME (Meng et al., 2022) could recall 99.70% editing facts in the editing direction, it can only recall 0.26% editing facts in the reverse direction.
@@
Towards bidirectional language model editing, a simple yet effective method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed to enhance the effectiveness of edited models in recalling the editing facts in the reverse direction. A set of editing objectives are designed to incorporate the bidirectional and invertible relationships between subject and object in a given fact into the updated model weights. Inspired by Mikolov et al. (2013), the representation of the newly editing object is driven close to the representation of the subject plus forward relation. Reversely, the representation of the subject is also driven close to the representation of the editing object plus the backward relation. Here, the relationships along both directions are inversible to each other. By this means, it can help characterize the bidirectional and inversible relationships between subject and object in factual knowledge, enabling models to establish factual associations bidirectionally during the editing process. Moreover, the proposed editing method eliminates the training step and does not require extra training data. Experiments with four representative LLMs of different sizes demonstrate that the proposed editing method not only improves the generalization ability when recalling the editing facts in the direction of editing, but also achieves highly promising performance in the reverse direction.
@@
In summary, our contributions are three-fold: (1) This study makes the first attempt to explore bidirectional language model editing and evaluation. A new evaluation metric of reversibility is introduced and a new benchmark BAKE is constructed to assess the reversibility of editing methods. (2) This paper presents meticulous experiments on existing popular editing methods and LLMs. Through analysis, we surprisingly found that existing methods and LLMs suffer serious deficiencies when evaluated in the reverse direction of editing. (3) A simple yet effective method BIRD is proposed to incorporate the bidirectional and inversible relationships between subject and object in a given fact into the updated model weights, which mitigate the reversal curse in model editing. To facilitate others to reproduce our results, we will publish all data and source code later at https://github.com/mjy1111/BAKE. We hope that both our benchmark and method can shed light on bidirectional language model editing.
@@ 2 Related Work
Many studies have investigated model editing, including memory-based, meta-learning, and locate-then-edit (Yao et al., 2023). Memory-based methods do not modify model weights but store the editing facts with an external memory (Mitchell et al., 2022b; Zhong et al., 2023). For example, Mitchell et al. (2022b) stored edits in a base model and learned to reason over them to adjust its predictions as needed. The latter two classes of methods are developed to directly modify the internal parameters of models, which is the focus of this paper. On the one hand, meta-learning methods train a hypernetwork to get gradient changes to update model parameters (Cao et al., 2021; Mitchell et al., 2022a). Cao et al. (2021) utilized a hypernetwork to predict parameter shift at test time. Mitchell et al. (2022a) learned to transform the fine-tuning gradient into a low-rank decomposition of the gradient. On the other hand, locate-then-edit methods first locate knowledge neurons in LLMs that exhibit a positive correlation with a knowledge expression, and then modify them accordingly (Dai et al., 2022; Meng et al., 2022, 2023). In particular, Dai et al. (2022) computed the contribution of each neurons to a certain knowledge, then updated or erased knowledge by modifying these neurons with the embedding vectors of facts. Meng et al. (2022) located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair in the MLP module.
@@
Recent works have shown a surprising failure of generalization in auto-regressive LLMs. Berglund et al. (2023) have verified that a model fine-tuned on a sentence of the form “A is B”, it will not automatically generalize to the reverse direction “B is A”. Moreover, the likelihood of the correct answer will not be higher than for a random answer. This is termed as the reversal curse. Contemporary to this work, Grosse et al. (2023) used influence functions to determine how much adding a given training example influences an LLM’s outputs. For instance, given the input A, what most influences the likelihood of B? In their experiments, training examples that match the order (“A precedes B”) are far more influential than examples with reverse order (“B precedes A”). Results show that if a pretrained model was not trained on facts in both directions, it would not generalize to both directions.
Interpretability of Transformer
Many studies have attempted to explain and understand Transformer (Vaswani et al., 2017). Previous works focus on the function of self-attention layers and the differences between layers (Tenney et al., 2019; Vig & Belinkov, 2019; Voita et al., 2019; Clark et al., 2019). For instance, Voita et al. (2019) found that the most confident attention heads played consistent linguistically-interpretable roles. More recently, a surge of works have investigated the role of feed-forward network (FFN) layers in Transformer, showing that FFN layers store factual and linguistic knowledge, which can be located in specific neurons and edited (Geva et al., 2021; Da et al., 2021; Meng et al., 2022; Dai et al., 2022). Besides, Geva et al. (2022) investigated the mechanism in which FFN layers update the inner representations of Transformer-based LMs. They propose that the FFN output can be viewed as a collection of updates that promote concrete concepts in the vocabulary space.
Compared with previous studies (Dai et al., 2022; Meng et al., 2022, 2023; Mitchell et al., 2022a; Yao et al., 2023; Zhong et al., 2023) that are the most relevant to our work, a main difference should be highlighted. These approaches target only on unidirectional editing and evaluation of model editing, while this study explores bidirectional editing and evaluation. To the best of our knowledge, this paper makes the first attempt to introduce a new evaluation metric of reversibility, build a benchmark for evaluating the reversibility of editing methods, and propose a simple yet effective method to mitigate the reversal curse in model editing.
@@ 3 Preliminary
This section introduces the definition of model editing and the preliminary knowledge of a popular editing method ROME (Meng et al., 2022).
3.1 Problem Definition
The goal of model editing is to insert new facts into model parameters without retraining. In this paper, we study facts of the form ($s,r,o)$, consisting of a subject $s$, a relation $r$, and an object $o$ (e.g., $s$ = Eiffel Tower, $r$ = locate in, $o$ = Paris). Editing a fact is to insert a new knowledge triple $(s,r,o^{*})$ in place of the current triple $(s,r,o)$, where these two triples share the same subject and relation. An editing operation is represented as $e=(s,r,o,o^{*})$ for brevity.
Given a set of editing facts =$\mathcal{E}=\left\{e_{1},e_{2},\ldots\right\}$ and a model $f$ , model editing involves learning a function $K$ that yields an edited language model $f^{*}:K(f,\mathcal{E})=f^{*}$. To evaluate the effectiveness of various model editing methods, previous works focus on evaluation along three dimensions: efficacy, generalization and locality (Cao et al., 2021; Mitchell et al., 2022a; Meng et al., 2022).
Efficacy
Given an editing fact ($s$ = Eiffel Tower, $r$ = located in, $o$ = Paris, $o^{*}$=London), it could be regarded edited effectively if the edited model $f^{*}$ assigns a higher probability to the statement “Eiffel Tower is located in London” than the original prediction (Paris).
Generalization
Edited models should be able to recall the updating knowledge when prompted within the editing scope. The paraphrase prompts of the question like “Where is the Eiffel Tower?” or “What is the location of the Eiffel Tower?” can be used. The edited model $f^{*}$ generalizes successfully if it could recall the editing memory “London”.
Locality
The edited model $f^{*}$ should remain unchanged on the irrelevant examples out of the editing scope. For example, the answer to “Where is the Louvre?” should still be “Paris”.

### v2

@@ Abstract
Large language models (LLMs) are prone to hallucinate unintended text due to false or outdated knowledge. Since retraining LLMs is resource intensive, there has been a growing interest in model editing. Despite the emergence of benchmarks and approaches, these unidirectional editing and evaluation have failed to explore the reversal curse. In this paper, we study bidirectional language model editing, aiming to provide rigorous evaluation to assess if edited LLMs can recall the editing knowledge bidirectionally. A metric of reversibility is introduced and a benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed to evaluate if post-edited models can recall the editing knowledge in the reverse direction of editing. We surprisingly observe that current editing methods and LLMs, while effective in recalling editing facts in the direction of editing, suffer serious deficiencies when evaluated in the reverse direction. To mitigate the reversal curse, a method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed, which designs a set of editing objectives to incorporate bidirectional relationships between subject and object into the updated model weights. Experiments show that BIRD improves the performance of four representative LLMs of different sizes via question answering and judgment.
@@
Figure 1: Comparison of unidirectional evaluation paradigms that assess whether edited models can recall the editing facts (a) via single-hop questions, or (b) via entailed questions in the direction of editing; and (c) the proposed bidirectional paradigm that assesses model editing in the reverse direction of editing. $f_{\theta}$ / $f_{\theta_{e}}$ denotes the models before / after editing.
@@ 1 Introduction
Large language models (LLMs) have shown impressive capabilities in understanding and generating text, improving performance on a variety of tasks (Chen et al., 2021; Ouyang et al., 2022; Zhang et al., 2023a). However, existing LLMs inevitably exhibit hallucinations due to false or outdated knowledge stored in their parameters (Zhang et al., 2023b; Peng et al., 2023; Ji et al., 2023). Considering the huge computing resource consumption of retraining LLMs and echoing green deep learning (Schwartz et al., 2020; Xu et al., 2021), there has been considerable attention on the research regarding knowledge editing (a.k.a., model editing(Sinitsin et al., 2020; Cao et al., 2021; Meng et al., 2022, 2023; Dai et al., 2022; Yao et al., 2023; Ma et al., 2024b), which aims at modifying knowledge by updating model weights.
@@
Existing methods developed to modify the parameters of models can be generally categorized into two paradigms of meta-learning or locate-then-edit (Wang et al., 2023; Yao et al., 2023). The effectiveness of an editing method is usually evaluated along three dimensions of efficacy, generalization and locality. The ZsRE (Levy et al., 2017) and CounterFact (Meng et al., 2022) datasets are primarily utilized to assess whether edited models can recall the newly injected facts, as depicted in Figure 1(a). Besides, Yao et al. (2023) and Zhong et al. (2023) have introduced portability and MQuAKE respectively to extend the assessment of generalization by measuring whether edited models can correctly answer the questions entailed by the editing facts, as shown in Figure 1(b). Despite the emergence of benchmarks and techniques for editing, they all operate under the unidirectional paradigm following only the direction being edited. The reversal curse (Berglund et al., 2023) has not yet been explored in model editing, and it remains unclear whether edited models can reason and recall the editing facts in the reverse direction of editing, as shown in Figure 1(c).
@@
In this paper, we study bidirectional language model editing, determining if edited LLMs can accurately recall the editing knowledge bidirectionally. Among various kinds of learned beliefs such as logical, spatial, or numerical knowledge, this paper investigates factual knowledge within LLMs, where each piece of knowledge is represented in the form of a (subject, relation, object) triple. A metric of reversibility is introduced to evaluate the effectiveness of edited models in recalling knowledge in the reverse direction of editing. A benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed. When inverting the relation between subject and object, the taxonomy which categorizes relations between entities into four classes: one-to-one, one-to-many, many-to-one, many-to-many (Bordes et al., 2013), is considered. To characterize these symmetrical and asymmetrical relations, two evaluation forms of question answering (QA) and judgment are adopted for evaluation respectively.
@@
Towards bidirectional model editing, a method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed to assist edited models to recall the editing facts in the reverse direction. A set of editing objectives is designed to incorporate the bidirectional and invertible relationships between subject and object into the updated model weights. Specifically, the representation of the newly editing object is driven close to the representation of the subject plus forward relation. Reversely, the representation of the subject is also driven close to the representation of the editing object plus the backward relation. Here, the relationships along both directions are inversible to each other. By this means, it can help characterize the bidirectional and inversible relationships between subject and object in factual knowledge, enabling models to establish factual associations bidirectionally during the editing process.
To establish the comparison of existing model editing methods on bidirectional model editing, this paper presents meticulous experiments involving as many methods as possible such as FT (Zhu et al., 2020), MEND (Mitchell et al., 2022), KN (Dai et al., 2022), ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023). Four representative LLMs of different sizes are selected as the base models including GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang & Komatsuzaki, 2021), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). We surprisingly observe that current editing methods and LLMs, while effective in recalling editing facts in the direction of editing, suffer serious deficiencies when evaluated in the reverse direction. Strikingly, the LLaMA-2 model edited by the state-of-the-art ROME (Meng et al., 2022) can recall 99.70% of the editing facts in the editing direction, but only 0.26% of the editing facts in the reverse direction. Besides, experiments demonstrate that the proposed editing method BIRD not only improves the generalization ability when recalling the editing facts in the direction of editing, but also achieves highly promising performance in the reverse direction.
@@
In summary, our contributions are three-fold: (1) This study makes the first attempt to explore bidirectional language model editing. A metric of reversibility is introduced and a benchmark of BAKE is constructed to assess the reversibility of edited models. (2) This paper conducts meticulous experiments and presents surprising finding that existing methods and LLMs suffer from the severe reversal curse on model editing. (3) The BIRD method is proposed to mitigate the reversal curse in model editing, which incorporates the bidirectional and inversible relationships between subject and object into the updated model weights. All data and code will be available to facilitate reproducing our results. We hope that both our benchmark and method can shed light on bidirectional language model editing.
@@ 2 Related Work
Many studies have investigated model editing to modify the parameters of models. On the one hand, meta-learning methods train a hypernetwork to get gradient changes to update parameters (Cao et al., 2021; Mitchell et al., 2022). Cao et al. (2021) utilized a hypernetwork to predict parameter shift at test time. Mitchell et al. (2022) learned to transform the fine-tuning gradient into a low-rank decomposition of the gradient. On the other hand, locate-then-edit methods first locate knowledge neurons in LLMs that exhibit a positive correlation with a knowledge expression, and then modify them accordingly. In particular, Dai et al. (2022) computed the contribution of each neurons to a certain knowledge, then updated or erased knowledge by modifying these neurons with the embedding vectors of facts. Meng et al. (2022) located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair (Geva et al., 2021) in the MLP module. Besides, some benchmarks are proposed for assessing knowledge editing (Meng et al., 2022; Cohen et al., 2023; Ma et al., 2024a; Li et al., 2023; Gandikota et al., 2023). For example, Zhong et al. (2023) used the multi-hop questions to assess editing, suggesting that editing a particular fact implies that other facts need to be changed.
@@
Recent works have shown a surprising failure of generalization in auto-regressive LLMs. Berglund et al. (2023) have verified that a model fine-tuned on a sentence of the form “A is B”, it will not automatically generalize to the reverse direction “B is A”. Moreover, the likelihood of the correct answer will not be higher than for a random answer. This is termed as the reversal curse. Contemporary to this work, Grosse et al. (2023) determined how much adding a given training example influences an LLM’s outputs. For instance, given the input A, what most influences the likelihood of B? In their experiments, training examples that match the order (“A precedes B”) are far more influential than examples with reverse order (“B precedes A”). Results show that if a pretrained model was not trained on facts in both directions, it would not generalize to both directions.
Compared with previous studies (Dai et al., 2022; Meng et al., 2022, 2023; Mitchell et al., 2022; Yao et al., 2023; Zhong et al., 2023) that are the most relevant to our work, a main difference should be highlighted. These approaches target only on unidirectional editing and evaluation of model editing, while this study explores bidirectional editing and evaluation. To the best of our knowledge, this paper makes the first attempt to introduce a metric of reversibility, build a benchmark for evaluating the reversibility of editing methods, and propose a simple yet effective method to mitigate the reversal curse in model editing.
@@ 3 Preliminary
3.1 Model Editing
The goal of model editing is to insert new facts into model parameters without retraining. In this paper, we study facts of the form ($s,r,o)$, consisting of a subject $s$, a relation $r$, and an object $o$ (e.g., $s$ = Eiffel Tower, $r$ = locate in, $o$ = Paris). Besides, we follow previous works (Zhong et al., 2023; Yao et al., 2023) to employ discrete prompts (question or cloze-style statement) to test whether a fact is in a language model. Editing a fact is to insert a new triple $(s,r,o^{*})$ in place of the current triple $(s,r,o)$, where these two triples share the same subject and relation. An editing operation is represented as $e=(s,r,o,o^{*})$ for brevity.
Given an edit $e$ and a model $f$ , model editing involves learning a function $K$ that yields an edited language model $f^{*}:K(f,e)=f^{*}$. To evaluate the effectiveness of various model editing methods, previous works focus on evaluation along three dimensions: efficacy, generalization and locality (Mitchell et al., 2022; Meng et al., 2022).
Efficacy validates whether the edited models could recall the exact editing fact when presented with the sole editing prompt $p$. The assessment is based on Efficacy Score (ES), which represents the new object $o^{*}$ exceeds $o$ in probability after model editing:  $\,\mathbbm{1}[P_{f^{*}}(o^{*}\,|\,p)>P_{f^{*}}(o\,|\,p)]\,$, where $\mathbbm{1}$ is the indicator function.
[…]
Locality [^1] Also referred to as specificity in Meng et al. (2022, 2023). is to verify whether the output of the edited models for inputs out of editing scope remains unchanged after editing under locality prompts. via Locality Score (LS): $\mathbb{E}_{p_{l}\in\mathcal{P}^{L}}[\mathbbm{1}[\,P_{f^{*}}(o_{l}\,|\,p_{l})>P_{f^{*}}(o^{*}\,|\,p_{l})]\,]$, where $o_{l}$ was the original answer of $p_{l}$.
For example, given an editing fact ($s$ = Eiffel Tower, $r$ = located in, $o$ = Paris, $o^{*}$=London), the editing prompt $p$ could be “Eiffel Tower is located in”, A paraphrase prompt could be “What is the location of the Eiffel Tower?”. A locality prompt and its original answer could be “Where is the Louvre?” and “Paris” respectively.